OpenStudy (anonymous):

List the first four terms of a geometric sequence with t1 = 4 and tn = -3tn-1.

OpenStudy (anonymous):

t2 = -3*t1 t1 = 4 so t2 = -3*4 t2 =-12 t3 = -3 * t2 t3 = 36 t4 = -3*t3 t4 = -108 so the first 4 terms are 4, -12, 36, and -108